An unknown gas at 77.1 

C


 and 1.05 atm

 has a molar mass of 44.10 g/mol.

 Assuming ideal behavior, what is the density of the gas?


1
Expert's answer
2021-02-26T06:05:09-0500

We have given,


Temperature, T = "77.1^{\circ} C" ="350.1K"


Pressure, P = "1.05" atm


Molar Mass, "M = 44.10 g\/mol"


Let the density of gas be d


Using the Ideal gas equation,


"PV = nRT"


"1.05\times V = \dfrac{m}{44.10}\times 8.314\times 350.1"


"\dfrac{m}{V} = \dfrac{46.305}{2910.73}"


"d = 0.0159g\/mol"